OpenStudy (anonymous):

How many liters of methane gas (CH4) need to be combusted to produce 12.4 liters of water vapor, if all measurements are taken at the same temperature and pressure? Show all of the work used to solve this problem. CH4 (g) + 2 O2 (g) yields CO2 (g) + 2 H2O (g)

OpenStudy (anonymous):

My Answer: n(H2O) = 12.4 / Vm n(CH4) = [(12.4 / 2) / Vm] = 6.2 / Vm V(CH4) = 6.4 / Vm * Vm = 6.2 L

OpenStudy (anonymous):

Is it correct?

OpenStudy (anonymous):

@e.mccormick @ganeshie8

OpenStudy (anonymous):

@SolomonZelman @.Sam.

OpenStudy (cuanchi):

Perfect!!!

OpenStudy (anonymous):

Thanks! :)